OpenStudy (anonymous):

Variable y varies directly with x2, and y = 72 when x = 6. Enter the constant of variation for the quadratic variation equation. y = ___x^2

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

72=k(6^2) solve for k, which is the same question as the last one you just posted lol

OpenStudy (anonymous):

36?

OpenStudy (dtan5457):

72=k(36)

OpenStudy (dtan5457):

k=36?

OpenStudy (anonymous):

So 36 is what I put in?

OpenStudy (dtan5457):

Does 36 times 36=72?

OpenStudy (anonymous):

Oh!

OpenStudy (dtan5457):

No...36 times_____=72?

OpenStudy (anonymous):

I forgot the last step. 2?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Yeah, it's 2.

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

can you check this out? http://brainly.com/question/264184 @dtan5457

OpenStudy (dtan5457):

y=kx^2 32=k(4^2) 32=k(16) what is k?

OpenStudy (anonymous):

2?

OpenStudy (dtan5457):

Yes so your function would be y=2x^2

OpenStudy (dtan5457):

just find which graph has that

OpenStudy (anonymous):

D?

OpenStudy (dtan5457):

Both pictures are the same thing lol

OpenStudy (dtan5457):

All I see is A and B

OpenStudy (anonymous):

Oh I didnt mean to do that one sec

OpenStudy (dtan5457):

It appears as if one of the answers are already there, but why not post just to make sure..

OpenStudy (anonymous):

Do you see it @dtan5457

OpenStudy (anonymous):

Please help It's my last question.

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

y=2x^2 Definitely not D.. One of d's points are 3,3 If you plug 3 into that function, you get 18 as y..

OpenStudy (anonymous):

b? @dtan5457

OpenStudy (anonymous):

a?